the sum of twice one number and three times a second number is 84 sum of thrice the first number and five times the second is 136.what is the number?

Answer by vidhyak(98)

Let the 2 numbers be x,y
the sum of twice one number and three times a second number is 84
2x + 3y =84 ......eq 1

sum of thrice the first number and five times the second is 136
3x +5y =136 ......eq 2

Multiply eq1 by 3 and eq 2 by 2 to cancel out 'x'
6x + 9y =252
6x +10y =272

Subtract the above equations
-y = -20
y = 20

Substitute y = 20 in eq 1
2x + 3y =84
2x + 3(20) =84
2x + 60 =84
2x =84 -60
2x = 24
x =12
x = 12, y = 20